Given that P(A) - o.5, P(B) = 0.3, and P)A and B)= 0.10, determine (A/B)?
answers:
0.125
0.20
0.33
0.50

Answer by stanbon(75887)

Given that P(A) - o.5, P(B) = 0.3, and P)A and B)= 0.10, determine (A/B)?
------------
P(A|B) = P(A and B)/P(B) = 0.1/0.3 = 1/3 = 0.33
-------------
Cheers,
Stan H.
==============
answers:
0.125
0.20
0.33
0.50